OpenStudy (anonymous):

In a survey of 300 members of a local sports club, 150 members indicated that they plan to attend the next Summer Olympic Games, 90 indicated that they plan to attend the next Winter Olympic Games, and 60 indicated that they plan to attend both games. How many members of the club plan to attend

OpenStudy (anonymous):

i think for the next Summer Olympic Games = 150 + 60 = 210, while next Winter Olympic Games = 90 + 60 = 150